# SCEP Collaboration The projected sensitivity of SCEP experiment to Magnetic Monopole

### v2

@@ V.4 Total Background and Sensitivity
Low-speed MMs () are unable to produce enough lights to surpass the energy threshold in PS. Searches of such MMs need to be performed with induction signal only, and higher number of fired coils is required. The final background rate can be expressed as:
@@
Here denotes as the radius of a single coil, and we set it to be 6 cm in the following calculation. The and are given in Subsec. V.2 and V.3, respectively. The induction threshold is optimized based on the sensitivity, which is equivalently the mean Feldman-Cousins upper limit [feldman1998unified] of the final coincidence background model under the background only hypothesis. The optimized threshold depends on the SNR0, the total exposure, the assumed MM speed and the coincidence number:
@@
where is the acceptance of PSs due to energy threshold the angular cut-off , is the acceptance of induction coils from Eq. 10, and is the geometrical acceptance of coil to MMs. is the assumed exposure. denotes as the Feldman-Cousins upper limit at CI under background only hypothesis, when events are observed with background predicted. is the Poisson probability of observing events under predicted background of .
![Refer to caption](x21.png)
@@
The estimated sensitivities of MM flux for moon-based detector and terrestrial detector are given in Fig. 18 and Fig. 19, respectively. We also plot the most stringent constraints for MM flux at different speed ranges from all the induction experiments [giacomelli1984magnetic], MACRO [macro2002final], and IceCube [aartsen2016searches]. With sufficient SNR of induction detection, sufficient number of induction coil layers , and the coincidence between the induction and scintillation signals, SCEP has the potential to achieve excellent background suppression for Dirac MMs traveling at speeds exceeding 2.510-4 light speed, in cases of both the terrestrial and moon-based detectors. Consequently, the sensitivities within this speed range are primarily dominated by the exposure (the product of exposure time and area of detection area). In cases of not sufficient or SNR, when the velocity of the MMs falls within the range of approximately 2.510-4 to 210-3, the scintillation background becomes significant, resulting in a reduction in sensitivity. This effect is particularly noticeable when number of coil layer Nc is less than 2, the SNR0 is less than 4, and when the detector is based on the Moon. For MMs traveling at speeds below approximately 2.510-4 light speed, the Dirac MMs are unable to produce scintillation lights in the PSs, causing SCEP to operate solely in induction-only search mode. As a result, sensitivities in this speed range decrease significantly due to the absence of scintillation/induction coincidence. However, the use of a higher number of coil layers can help recover the lost sensitivity to some extent. Also, the sensitivity keeps decreasing as the speed of MM decreases due to the SNR’s dependence on the MM speed.
@@ VI Summary and Discussion
The SCEP experiment aims to detect the induction signal and scintillation signal simultaneously when a MM passes through coils and deposits energy inside PSs. Two read-out scenarios are planned for the induction signal. It can either be directly amplified by an operational amplifier and read out by a digitizer, or the induction micro-current can be fed into a Helmholtz coil, converting it to a magnetic signal that can be read out by a high-precision magnetometer coupled to the Helmholtz coil. The background for the induction signal is primarily influenced by two factors: the backend electronics noise and the thermal noise from the induction coil. In the case of the magnetometer-readout scenario, the backend noise is negligible, and the background is dominated by the thermal noise. To estimate the SNR for three prototype induction coils, a dedicated simulation framework has been developed. The results show that the V2 coil with magnetometer readout can achieve an SNR of approximately 2. These findings are further validated through measurements using stimulated pulsed magnetic flux from a long-thin coil and measurements of the thermal noise on the coil. The pileup background for the scintillation signal on the PSs is estimated using dedicated MC simulations, taking into account the particle fluxes at the detector site in both terrestrial (sea-level) and moon-based environments. The pileups are primarily caused by muons and protons for terrestrial detectors, while high-energy protons and helium nuclei dominate in moon-based detectors. By considering all these factors, the sensitivities of the SCEP experiment to Dirac MMs are estimated as a function of MM speed, with different assumed exposures. With sufficiently high SNR on each coil, it is promising to reach a background-free search of Dirac MM with the help of PS coincidence, even in a moon-based detector with high flux of cosmic rays. However, for MMs traveling at speeds below approximately 2.510-4 light speed, they are unable to produce sufficient scintillation light in the PSs. Consequently, induction-only searches have to be performed in this speed range, and the sensitivity is reduced unless a sufficient number of coil layers are used. With Nc=3, SNR0¿4, and an exposure of 500 myr, the SCEP experiment can already achieve the best sensitivity for MM searches at speeds around 10-5 light speed in terrestrial environment.
However, it should be noted that the estimations presented in this work are based on a coil size with a radius of 6 cm. In future large-area detector arrays, there is a preference for larger coil sizes and smaller number of coil layers due to practical considerations such as reducing the number of readout channels and the total weight of the system. In addition, we currently achieved an SNR02 for such coil. Further optimization is required to reach SNR4. In the ideal case, the SNR on a single coil is proportional to the product of the number of turns, the square of the wire diameter, and the inverse square of the coil diameter, and thus decreases if we simply increase the coil size. This relationship does not even take into account the potential dependence of the alternating resistance on these parameters. On the other hand, increasing the coil size will lead to a decrease in the track reconstruction resolution when using induction signals from multiple coils. Consequently, this can result in a higher background rate after the induction/scintillation coincidence. Therefore, when employing detector arrays with larger coil sizes, it is crucial to optimize the SNR for the induction signal and minimize the background rate of the scintillation signal on the PSs. To improve the SNR of the induction signal, optimization of the induction circuit and the use of materials with high magnetic permeability are two potential approaches. Increasing the number of layers in the particle detectors and incorporating different types of particle detectors hold promise for further reducing the background rate due to particle interactions. Additionally, more advanced algorithms, such as those based on deep neural networks, have the potential to enhance background rejection by leveraging the full range of information obtained from the data.
Taking all of these factors into consideration, it is promising that SCEP is capable of conducting background-free searches for medium-to-high-speed Dirac magnetic monopoles by requiring the coincidence between the induction and scintillation signals. It is also possible to further increase the SNR of induction searches, allowing for the use of fewer coil layers and increased sensitivity for low-speed magnetic monopoles.

### v3

@@ V.4 Total Background and Sensitivity
Low-speed MMs () are unable to produce enough lights to surpass the energy threshold in PS. Searches of such MMs need to be performed with induction signal only. The total background rate per unit area can be expressed as:
@@
Here denotes as the radius of a single coil, and we set it to be 6 cm in the following calculation. The and are given in Subsec. V.2 and V.3, respectively. The induction threshold is optimized based on the sensitivity, which is equivalently the mean Feldman-Cousins upper limit [feldman1998unified] under the background-only hypothesis. The optimized threshold depends on the SNR0, the total exposure, the assumed MM speed and the coincidence number:
@@
where is the acceptance of PSs due to energy threshold and the angular cut-off . is the trigger acceptance of the induction coils from Eq. 10, and is the geometric acceptance of the induction coils to the MMs. is the assumed exposure. denotes as the Feldman-Cousins upper limit at CI under background only hypothesis, when events are observed with background predicted. is the Poisson probability of observing events under predicted background of .
@@
The estimated 90% exclusion sensitivities of the MM flux as a function of the MM speed for a terrestrial detector are given in Fig. 17, with different assumptions of the SNR0, Nc, and exposure. And we also give the sensitivities for a moon-based detector in the appendix. The most stringent constraints to the MM flux at different speed ranges from all the induction experiments [giacomelli1984magnetic], MACRO [macro2002final], and IceCube [iceCube2022_mm_result, abbasi2022search], as well as the astrophysical constraints (Parker boudary [parker1970origin] and extended Parker boundary [adams1993extension, lewis2000protogalactic]), are also plotted for comparison. SCEP has the potential to achieve excellent background suppression for cosmic Dirac MMs traveling at speeds exceeding 2.510-4 light speed. The sensitivities within this speed range are primarily dominated by the exposure (the product of exposure time and area of detection area). For MMs traveling at speeds below approximately 2.510-4 light speed, the Dirac MMs are unable to produce scintillation lights in the PSs, causing SCEP to operate solely in induction-only search mode. As a result, sensitivities in this speed range decrease due to the absence of the scintillation/induction coincidence. However, the use of a higher number of coil layers and a larger value of SNR can help recover the lost sensitivity to some extent. We estimate that a SNR and a N are required to reach the current best limits by all the induction searches [giacomelli1984magnetic] with an exposure of 20000 yearm2. Also, the sensitivity keeps decreasing as the speed of MM decreases due to the SNR’s dependence on the MM speed.
V.5 Acceptance Loss due to Earth Shielding
The Earth’s matter can shield the cosmic MMs with insufficient mass, causing an additional acceptance loss for a terrestrial MM detector. We perform a simple estimation of the acceptance loss due to this Earth shielding effect, assuming the Earth is a sphere composed of only the atmosphere, silicon mantle, and iron core layers. The thicknesses of these three layers are taken as 2000, 3500, and 2900 km, respectively. The air density distribution along the altitude uses the data from Ref. [Characterisations]. The energy losses of the MMs in the air, the silicon, and iron follow the model given in Ref [DERKAOUI1998173]. Particularly, the energy losses in the air for the MM in 10-3 to 10-2 and 10-5 to 10-4 light speed are scaled from that of the a proton and of the liquid helium, respectively, same as in Ref [DERKAOUI1998173]. We also neglect the scattering of the MMs inside the Earth and the atmosphere. The acceptance loss of the isotropic cosmic MMs due to the Earth shielding effect for a terrestrial sea-level detector is calculated and shown in Fig. 18.
For the upgoing cosmic MMs, they are not able to penetrate Earth’s mantle and core if their masses are less than 1011 to 1015 GeV (depending on the MM speed), leading to the reduction of the acceptance by about half. When the MM mass is less than certain limit (for example, 107 GeV for of 0.001), no cosmic MMs can be searched for using sea-level detectors since the atmosphere absorbs all of the downgoing MMs.
@@ VI Summary and Discussion
The SCEP experiment aims to detect the induction signal and scintillation signal simultaneously when a cosmic MM passes through. This dual-signal detection approach not only can provide smoking-gun signature for a potential cosmic MM traversal, but also suppresses the background caused by the environmental charged particles in the PSs and thermal noise in the induction coils. This allows the experiment to be deployed in a room-temperature environment at sea-level (or even high-altitude) locations, which is more cost-effective compared to experiments requiring an underground environment or cryogenic detector conditions. The SCEP design holds the potential to achieve a large-area detector array, which is crucial in reaching the MM flux sensitivity beyond the current experimental and astrophysical constraints.
We have developed a dedicated simulation framework for estimating the induction signal performance. This simulation has been validated through tests performed on three prototype induction coils with distinct geometric parameters, using both the magnetometer readout and the direct ADC readout. The differences between the model predictions and measurements in terms of signal amplitudes, signal shapes, and noise amplitudes are within 10%. The thermal noise of the induction coil, which increases with frequency due to the alternating resistance, is the main contributor of the induction background. Based on the simulation of the induction signal and an estimation of the environmental particle backgrounds, we provide an assessment of the sensitivity to the cosmic MM flux. Thanks to the coincidence between the induction and scintillation signals, the SCEP experiment is able to achieve a “background-free” search mode for MM speeds greater than approximately 2.510-4 the light speed. In this speed range, the background is significantly suppressed, and the sensitivity to the MMs is mostly proportional to the total exposure. An exposure of about 20,000 yearm2 will be sufficient to reach the current flux limit set by the MACRO experiment [macro2002final] over a wide range of the MM speeds (from 2.510-4 to 0.75 light speed). For MMs with speeds less than about 2.510-4 light speed, they are no longer able to create scintillation signals in the PSs. The SCEP experiment will operate in an induction-only mode. To reach the limits set by all previous induction-only experiments combined [giacomelli1984magnetic], an SNR0 greater than 4.5 and a number of coil layers more than 2 are preferred, given the assumed exposure of 20,000 yearm2.
The current three test coils, without optimization, have achieved an SRN02. Further optimization is required to reach the SNR0 larger than 4.5. To improve the SNR of the induction signal, two approaches are planned: 1) optimization of the induction circuit and 2) the use of materials with high magnetic permeability. The latter is of particular potential. A preliminary estimation shows that a high-permeability magnetic core can effectively amplify the MM signal by at least one order of magnitude, with the amplification factor depending on the core material and geometric dimensions. This amplification can be achieved without introducing significant additional thermal noise from the eddy currents and hysteresis within the frequency range of interest. This would result in an overall amplification of the SNR by several hundred times. Such a level of SNR is sufficient to meet the requirements for the planned large-area array detectors. More details on the optimization and performance improvements will be provided in future dedicated publications. Similar analyses have also been performed in Ref. [tumanski2007induction, coillot2015magnetic].
In addition to the induction signal optimization, increasing the number of the PS detector layers and incorporating different types of particle detectors hold promise for further reducing the background rate due to the environmental particles. Moreover, more advanced algorithms, such as those based on the deep neural networks, have the potential to enhance the background rejection by leveraging the full range of information obtained from the experimental data.
@@ Acknowledgements
Appendix A Sensitivity for a Moon-based Detector
A moon-based deployment of the SCEP detector is of particular scientific interest, not only for the cosmic MM searches but also for the detection of other exotic particle phenomena. The magnetometer of the SCEP detector can be specialized to search for the pseudo-magnetic fields that could be induced by the presence of exotic particles [jiang2021search]. A super-long baseline detector network, with magnetometers located on both the Earth and the Moon, can potentially increase the sensitivity for the detection of the topological defect dark matters, such as domain walls [lohe1979soliton, gani2014kink, gani2015kink]. Similar Earth-based detector network concepts have been proposed and implemented in previous works [afach2021search]. For MM searches specifically, a moon-based detector can extend the mass range of the search, as the Moon has no atmosphere and less intervening matter compared to a terrestrial detector. This can lower the detectable MM mass threshold by approximately 2 to 4 orders of magnitude. Fig. 19 shows the minimal detectable MM mass as a function of MM speed for both a moon-based and a terrestrial sea-level detector. In the calculation of the minimal detectable MM mass for a moon-based detector, it is assumed that the MM must be able to penetrate a 5-cm thick PS layer to be registered by the detector.
However, the particle background environment on the lunar surface is more severe compared to the terrestrial environment. On the Moon, muons are no longer the dominant source of background, due to the absence of the atmosphere. In the deep space environment surrounding the Moon, high-energy protons and helium nuclei emerge as the prevailing sources of the background particles, as shown by AMS [aguilar2021periodicities, aguilar2022properties], DAMPE [dampe2019measurement, dampe2022detection], and CALET [marrocchesi2019measurement, brogi2022measurement] experiments. Assuming the negligible influence of the Earth’s magnetic field on the lunar surface, it is conservatively assumed that the fluxes of protons, helium nuclei, carbon nuclei, and iron nuclei are approximately isotropic. The fluxes of these background particle species as a function of their kinetic energies are taken from Ref. [waller2020simulated], and are also shown in Fig. 20.
The estimated particle background for a moon-based detector is shown in Fig. 21. The calculation is based on the methods described in Subsection V.3. The dominant background is due to the cosmic protons, helium nuclei, carbon nuclei, and iron nuclei. These high-energy particles cover the entire energy deposition spectrum up to the GeV range, overlapping with the expected energy distributions for the MM of different speeds. As a result, the particle background for a moon-based detector is approximately 4 orders of magnitude higher than that of a sea-level terrestrial detector. The sensitivity of a moon-based MM detector is given in Fig. 22, based on the calculations outlined in Subsection V.4. Due to the high particle background in the MM speed range greater than 2.510-4 light speed, a higher trigger threshold of the induction signal is required to achieve a “background-free” search. This, in turn, lowers the overall acceptance of the detector. Consequently, the sensitivity of a moon-based MM detector is not as competitive as that of a sea-level terrestrial detector. The required SNR0 and Nc are raised to be larger than 7.6 and 2, respectively, to overcome the elevated particle background environment on the lunar surface, with an assumed exposure of 20000 yearm2.